OpenStudy (anonymous):

5x-y=-9 x-y=-1 can anyone help me solve this equation. here are my choices for this question. (1,2) (2,1) (-1,-2) (-2,-1)

OpenStudy (anonymous):

Looks like a system of equations that can be solved by elimination or substitution.

OpenStudy (anonymous):

can u help me with this problem

OpenStudy (anonymous):

Only if you are willing to show some effort.

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

Do you have any idea how you might start to solve this?

OpenStudy (anonymous):

honestly no first time doing this problem

OpenStudy (anonymous):

OK. Two options that I can help you with are: Substitution - solve one equation for one variable and substitute that into the other equation. Elimination - using addition or subtraction to eliminate one variable. They are usually taught at two different methods but I see them as variations on a theme.

OpenStudy (anonymous):

Elimination might be the most efficient way to solve this system of equations.

OpenStudy (anonymous):

\(5x-y=-9\) \(x-y=-1 \) The idea is to eliminate one of the variables. Any thoughts?

OpenStudy (anonymous):

5x-y=-10??

OpenStudy (anonymous):

I like the way you think, but it's not quite what I meant. Remember that you can do anything you want to an equation as long as you do the same thing to every term in the equation.

OpenStudy (anonymous):

Also remember that you can combine like terms.

OpenStudy (anonymous):

ohhh... ok makes a little bit more sense

OpenStudy (anonymous):

This will be a bit of the scenic route but I think it will help you to understand the process better. Try adding the two equations together and see what you get.

OpenStudy (anonymous):

how would i do that again

OpenStudy (anonymous):

Combine like terms: 5x + x -y + -y -9 +-1 Keep them in the same columns

OpenStudy (anonymous):

i dont get how i can set this up

OpenStudy (anonymous):

5x -y = -9 x -y = -1 ---------- 6x -2y = -10

OpenStudy (anonymous):

Do you see what I did?

OpenStudy (anonymous):

i kinda did

OpenStudy (anonymous):

do i sole that equation

OpenStudy (anonymous):

No, I did warn you that this was the long way around. You can do what ever you want to an equation as long as you do the same thing to each term in the equation. Using what I just showed you about adding the equations, what can you do to one of the equations to eliminate one of the variables?

OpenStudy (anonymous):

divide it by 2 o divide it by 6 on both sides

OpenStudy (anonymous):

That will reduce a variable to a coefficient of 1. 5x-y=-9 x-y=-1 Look at only these equations and try again. When you add them together as I did, you want one of the variables to add to 0.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Do you know what you need to do?

OpenStudy (anonymous):

no.. :(

OpenStudy (anonymous):

You need to change one equation so that adding the coefficients together will equal 0. Think additive inverse.

OpenStudy (anonymous):

1-1 = 0 -4+4=0 ...

OpenStudy (anonymous):

Look at the y terms. What can you multiply by to give them opposite signs?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\(-y \times 2 = -2y\)

OpenStudy (anonymous):

\(-y \times\) what? \(= 0\)

OpenStudy (anonymous):

4y

OpenStudy (anonymous):

-y + y = 0 Sorry, I was using the wrong operation. You want to add the equations so I should have said: \(-y + \) what? \(=0\)

OpenStudy (anonymous):

So, what can you multiply the -y by to change the sign to positive?